OpenStudy (anonymous):

Given the quadratic equation, written in General/ vertex form, state/ discuss.. Y=3/4(x+2)^2-5 1) the vertex 2) opening up or down 3) wider or narrower comparing with y=x^2 4)Domain and range 5) axis of symmetry PLEASE help!! this is a pre-test question I don't understand

OpenStudy (anonymous):

@whpalmer4 can you help?

OpenStudy (anonymous):

@kropot72 ? can you help

OpenStudy (anonymous):

@Hero

hero (hero):

If given a quadratic equation of the form y = a(x - h) + k, then (h,k) is the vertex \(a\) is positive, then the graph opens up \(a\) is negative then the graph opens down If \(0 < a < 1\) then the graph will be wider than y = x^2 If \(a > 1\) or \(a < -1\) then the graph will be narrower The domain of a quadratic equation is always all real numbers If the graph opens up then the range will be \(-b/(2a), \infty\) If the graph opens down, then the range will be \(-b/(2a), -\infty\)

OpenStudy (anonymous):

Thank you so much :D @Hero

hero (hero):

Well, you still have to do the work. I was just giving you general info.